# Knowledge Graph Retrieval-Augmented Generation for LLM-based Recommendation

### v1

@@ 1 Introduction
Recommender systems, as techniques designed to assist people in making decisions in their daily lives, are increasingly gaining impact in various fields (Kenthapadi et al., 2017; He et al., 2020; Fan et al., 2019), such as online shopping, job matching, and social media. Typically, recommender systems leverage user-item interactions, item characteristics, and user profiles to predict user preferences and provide relevant recommendations (He et al., 2020; Sarwar et al., 2001; Wang et al., 2024). Recently, Large Language Models (LLMs) have achieved significant breakthroughs, which further drive developments in various domains (Fan et al., 2024b; Zhao et al., 2024; Wu et al., 2023a). Especially with the success of LLMs, recommender systems have seen rapid growth (Geng et al., 2022; Bao et al., 2023; Qu et al., 2024). By training on a wide range of data, LLMs (e.g., GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023)) are able to acquire extensive knowledge and demonstrate exceptional language understanding capability. This capability enables LLM-based recommender systems to capture user preferences through a nuanced understanding of relevant attributes (e.g., user profiles, item descriptions, historical interactions) for more accurate recommendations. As a result, LLM-based recommender systems have emerged as a new paradigm in recommendation technology (Zhao et al., 2024).
However, despite their powerful language understanding and generalization capability, LLM-based recommender systems face significant challenges, including hallucinations and lack of up-to-date and domain-specific knowledge (Luo et al., 2023). Specifically, one key issue is that LLM-based recommender systems may generate recommendations that are entirely fictional due to the inherent limitations of LLMs. For example, an LLM-based recommender system may recommend ”Godmother”, a non-existent film, to the user who has watched ”The Godfather”, as illustrated in Figure 1 (a). Additionally, LLMs usually lack up-to-date knowledge, which can prevent them from recommending the latest films or products in a timely manner. As illustrated in Figure 1 (b), the LLM-based recommender system is unable to recommend the latest films due to the training data only containing up to December 2022. Furthermore, LLMs often lack domain-specific knowledge, as recommendation-oriented corpora are very limited during the training phase of LLMs (Geng et al., 2022). Consequently, LLMs may struggle to meet the nuanced needs of recommendation tasks. To alleviate these issues, one potential solution is to frequently fine-tune the LLMs with up-to-date and domain-specific knowledge. However, the massive parameters of LLMs make this process computationally expensive and time-consuming, which severely hinders the practical application in the real world.
@@
To address the aforementioned challenges, in this paper, we propose a knowledge retrieval-augmented recommendation framework, namely K-RagRec, to provide up-to-date and reliable knowledge by retrieving relevant knowledge from item’s KGs for recommendation generation in the era of LLMs. Specifically, our proposed framework first performs knowledge sub-graph indexing on the items’ KG at a coarse and fine granularity to construct the knowledge vector database. Next, an adaptive retrieval policy is designed to determine which items should be retrieved, followed by the retrieval of specific sub-graphs from the knowledge vector database. To refine the quality of retrieval and ensure the most relevant results are prioritized at the top of the input, we subsequently re-rank the retrieved knowledge sub-graphs. Finally, we introduce a GNN and projector to align the retrieved knowledge sub-graphs into the semantic space of the LLM for knowledge-augmented recommendation. The main contributions of this paper can be summarized as follows:
  • We propose a novel framework (K-RagRec) that retrieves faithful knowledge from KGs to augment the recommendation capability of LLM. Note that we introduce a flexible indexing method for KGs, which can provide a comprehensive view of a node’s neighborhood within KG.
  • We design an adaptive retrieval strategy to determine whether an item needs to be retrieved based on its popularity, significantly improving efficiency.
@@ 2 Related Work
In this section, we briefly overview some related studies focusing on RAG and LLM-based recommender systems.
2.1 Retrieval-Augmented Generation
Recently, RAG has emerged as one of the most representative technologies in the field of generative AI, combining the strengths of retrieval systems and language models (LM) to produce coherent and informative text. Early methods, such as REALM Guu et al. (2020), RETRO Borgeaud et al. (2022), and DPR Karpukhin et al. (2020), typically involve retrieving relevant fragments from a large corpus to guide the LM generation. However, standard RAG methods often struggle to accurately retrieve all relevant textual chunks, due to unnecessary noise and even harmful disturbance in the documents. To address these limitations, recent studies Baek et al. (2023); Wu et al. (2023b); He et al. (2024); Luo et al. (2023); Wang et al. (2023); Sen et al. (2023); Sun et al. (2023) focus on retrieving structured and faithful knowledge from graphs for enhancing generations. For example, Retrieve-Rewrite-Answer Wu et al. (2023b) retrieves relevant sub-graphs from KG and converts retrieved sub-graphs into text for the generation. G-Retriever He et al. (2024) explores retrieving sub-graphs from various types of graphs to alleviate the hallucinations of LLM.
With the explosion of LLMs in recommendations, a few works Di Palma (2023); Wu et al. (2024); Ang et al. (2024) make initial explorations in RAG for recommendations. For instance, the work Di Palma (2023) proposes leveraging knowledge from movie or book datasets to enhance recommendations. Nevertheless, retrieving faithful structured knowledge from the KGs for recommendations is under-explored and shows great potential. To fill this gap, we propose to retrieve knowledge sub-graphs from KGs to enhance the recommendation performance of LLM.
2.2 LLM-based Recommendation
The remarkable breakthroughs in LLMs have led to their widespread adoption across various fields, particularly in the recommendations. Given powerful reasoning and generalization capabilities, many studies have actively attempted to harness the power of the LLM to enhance recommender systems (Geng et al., 2022; Bao et al., 2023; Qu et al., 2024; Wei et al., 2024; Zhang et al., 2023; Ning et al., 2024). For example, P5 (Geng et al., 2022) proposes an LLM-based recommendation model by unifying pre-training, prompting, and prediction for various recommendation tasks, such as sequential recommendation and rating predictions. Furthermore, Tallrec (Bao et al., 2023) fine-tunes LLM (i.e., LLaMA-7B) to align with recommendation data for sequential recommendations. To further capture higher-order collaborative knowledge and enhance the model’s ability to generalize users and items, TokenRec (Qu et al., 2024) proposes a masked vector-quantized tokenizer to tokenize users and items in LLM-based recommendations. Despite their effectiveness, these models often face challenges, such as hallucinations and the lack of up-to-date knowledge. While fine-tuning can partially mitigate these issues, it is resource-costly and time-consuming due to the massive parameters of LLMs. To overcome these challenges, we propose a knowledge retrieval augmented recommendation framework that leverages external KGs to provide reliable and up-to-date knowledge instead of costly fine-tuning.
@@ 3.1 Preliminary
Knowledge Graph: In this work, we propose to leverage external knowledge databases (i.e., KGs) to augment LLMs for recommendations. KGs contain abundant factual knowledge in the form of a set of triples: $\mathcal{G}=\{(n,e,n^{\prime})\ |\ n,n^{\prime}\in N\ ,\ e\in E\}$, where $N$ and $E$ denote the set of entities and relations, respectively. For example, a triple $Interstellar\ \xrightarrow{directed\_by}\ Nolan$ indicates that the movie ”Interstellar” was directed by the director ”Nolan”.
Graph Neural Networks: GNNs are a critical technique in graph machine learning and are widely employed in various graph tasks. By iteratively updating node representations through aggregating information from neighboring nodes, GNNs effectively capture the underlying topology and relational structure of graphs. Formally, a typical GNN operation can be formulated as follows:
LLM-based Recommendations: Let $U=\{u_{1},u_{2},\dots,u_{n}\}$ and $V=\{v_{1},v_{2},\dots,v_{m}\}$ represent the sets of users and items, where $n$ and $m$ are the sizes of users and items, respectively. The goal of an LLM-based recommender system is to understand users’ preferences by modeling users’ historical items interactions $V^{u_{i}}=[v^{u_{i}}_{1},v^{u_{i}}_{2},\cdots,v^{u_{i}}_{|L_{u_{i}}|}]$ (e.g., clicks, bought, etc.), where $|L_{u_{i}}|$ is interaction sequence length for user $u_{i}$. Notably, item $v_{j}$’ side information, such as title and description, is publicly available to enhance LLM for modeling user preferences. In our setting, we consider asking the LLM to select user $u$’s preferred item $v$ from the candidate set $C=\{v_{j}\}^{M}_{j=1}$, where $M$ is the number of candidate items. The candidate set $C$ typically consists of one positive sample as well as $M-1$ negative samples. Specifically, for a frozen LLM $f_{\delta}$ with parameters $\delta$, we denote an input-output sequence pair as $(Q,A)$, where $Q$ is the recommendation query/prompt, which consists of task descriptions and users’ historical items. The output $A$ is the LLM’s prediction. Furthermore, we provide more details about the prompt template of LLM in appendix A.3.
Figure 2: The overview of the K-RagRec. It contains five key components: Hop-Field Knowledge Sub-graphs for Semantic Indexing, Adaptive Retrieval Policy, Knowledge Sub-graphs Retrieval, Knowledge Sub-graphs Re-Ranking, and Knowledge-augmented Recommendation.
@@ 3.2 The Overview of Proposed Method
As shown in Figure 2, our proposed K-RagRec consists of five crucial components: Hop-Field Knowledge Sub-graphs for Semantic Indexing, Adaptive Retrieval Policy, Knowledge Sub-graphs Retrieval, Knowledge Sub-graphs Re-Ranking, and Knowledge-augmented Recommendation. The model first performs indexing of hop-field knowledge sub-graphs within the KG. Following this, an adaptive retrieval policy is implemented to determine which items should be retrieved or augmented. The model then retrieves specific sub-graphs from the knowledge vector database. Subsequently, the retrieved knowledge sub-graphs are re-ranked to refine the quality of the retrieval. Finally, the retrieved knowledge sub-graphs are utilized with the original prompt to generate recommendations.
@@ 3.3 Hop-Field Knowledge Sub-graphs for Semantic Indexing
Typically, retrieving knowledge from the KG chunks the KG into nodes (He et al., 2024) or triplets (Luo et al., 2023) and only retrieves content locally around the target entity in the KG. However, these methods just naively retrieve the first-order neighbors of an entity (i.e., item), which makes it difficult to capture these higher-order neighborhood effects among entities/items in the recommendation process. Therefore, to effectively retrieve knowledge from the KG, we propose performing semantic indexing on hop-field knowledge sub-graphs, which can flexibly chunk KGs and provide a comprehensive view of a node’s neighborhood in KG. As illustrated in Figure 2 component 1, we first introduce a pre-trained language model (PLM), such as SentenseBert (Reimers, 2019), to capture the semantic information for node $n_{o}$ as follows:
@@
$\textbf{z}_{n_{o}}^{l}=\text{GNN}_{\phi_{1}}^{\text{Indexing}}(\{\textbf{z}_{n_{m}}^{(l-1)},\textbf{r}_{e_{<o,m>}}^{(l-1)}:~{}n_{m}\in\mathcal{N}(n_{o})\}),$ (4)
@@
3.4 Adaptive Retrieval Policy
Although RAG can augment the LLM for modeling user preferences with retrieved knowledge, retrieving each item can cost a significant amount of retrieval time, which can severely degrade the user experience and cause user churn. Meanwhile, due to abundant interactions between users and popular items, most models tend to keep rich knowledge of popular items, resulting in an inferior performance for cold-start items (Zhao et al., 2023). To this end, we design an adaptive retrieval policy to determine whether an item needs to be augmented from KG based on its popularity (e.g., sales volume and page view). Particularly, the item is retrieved if its popularity is less than the pre-defined threshold $p$, otherwise not. By incorporating this strategy, the retrieval time in K-RagRec can be significantly reduced to achieve more efficient retrieval.
@@ 3.5 Knowledge Sub-graphs Retrieval
Next, we retrieve top-$K$ most similar sub-graphs $G_{j}=\{g^{\prime}_{1},...,g^{\prime}_{K}\}$ from $\mathcal{G}$ for item $v_{j}$:
@@ 3.6 Knowledge Sub-graphs Re-Ranking
To refine the quality of retrieval for enhancing recommendation performance, the next crucial step is knowledge sub-graphs re-ranking. Feeding all retrieved knowledge sub-graphs $g^{\prime}\in G$ directly into LLM can lead to information overload if the user ${u_{i}}$ has a long historical interaction list towards items. Therefore, we execute re-ranking to shorten the retrieved knowledge sub-graphs and ensure the most relevant knowledge sub-graphs are prioritized at the top of the prompt. Specifically, we adopt the recommendation prompt as a query for re-ranking. For example, this recommendation prompt can be ”What is the top recommended movie for the user who watched {Matrix, …, Iron Man}?”. For consistency, we adopt the same PLM to capture the semantic information of the above prompt as p, and re-rank the knowledge sub-graphs in $G$ to obtain a Top-$N$ knowledge sub-graphs set $\hat{G}$:
@@ 3.8 Optimization for K-RagRec
Therefore, instead of fine-tuning the LLM model extensively, we only need to learn parameters of two GNNs (i.e., $\text{GNN}_{\phi_{1}}^{\text{Indexing}}$, $\text{GNN}_{\phi_{2}}^{\text{Encoding}}$) and projector $\text{MLP}_{\theta}$, while parameters $\delta$ of LLM backbone are frozen. We update the parameters $\phi_{1}$, $\phi_{2}$ and $\theta$ through the Cross-Entropy Loss $\mathcal{L}(Y,A)$, where $Y$ is the ground-truth and $A$ is LLM’s predition.

### v2

@@ Abstract
Knowledge Graph Retrieval-Augmented Generation for LLM-based Recommendation
Shijie Wang1, Wenqi Fan1thanks: Corresponding authors, Yue Feng211footnotemark: 1, Shanru Lin3, Xinyu Ma4, Shuaiqiang Wang4, Dawei Yin4, 1The Hong Kong Polytechnic University, 2University of Birmingham, 3City University of Hong Kong, 4Baidu Inc, shijie.wang@connect.polyu.hk; wenqifan03@gmail.com; y.feng.6@bham.ac.uk; lllam32316@gmail.com; xinyuma2016@gmail.com; shqiang.wang@gmail.com; yindawei@acm.org
@@ 1 Introduction
Recommender systems, as techniques designed to assist people in making decisions in their daily lives, are increasingly gaining impact in various fields (Kenthapadi et al., 2017; He et al., 2020; Fan et al., 2019), such as online shopping, job matching, and social media. Recently, Large Language Models (LLMs) have achieved significant breakthroughs, which further drive developments in various domains (Fan et al., 2024b; Zhao et al., 2024; Wu et al., 2023a). Especially with the success of LLMs, recommender systems have seen rapid growth (Geng et al., 2022; Bao et al., 2023; Qu et al., 2024). By training on a wide range of data, LLMs (e.g., GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023)) are able to acquire extensive knowledge and demonstrate exceptional language understanding capability. This capability enables LLM-based recommender systems to capture user preferences through a nuanced understanding of relevant attributes (e.g., user profiles, item descriptions, historical interactions) for more accurate recommendations. As a result, LLM-based recommender systems have emerged as a new paradigm in recommendation technology (Zhao et al., 2024).
However, despite their powerful language understanding and generalization capability, LLM-based recommender systems face significant challenges, including hallucinations and lack of up-to-date and domain-specific knowledge (Luo et al., 2023). Specifically, one key issue is that LLM-based recommender systems may generate recommendations that are entirely fictional due to the inherent limitations of LLMs. For example, an LLM-based recommender system may recommend "Godmother", a non-existent film, to the user who has watched "The Godfather", as illustrated in Figure 1 (a). Additionally, LLMs usually lack up-to-date knowledge, which can prevent them from recommending the latest films or products in a timely manner. As illustrated in Figure 1 (b), the LLM-based recommender system is unable to recommend the latest films due to the training data only containing up to December 2022. Furthermore, LLMs often lack domain-specific knowledge, as recommendation-oriented corpora are very limited during the training phase of LLMs (Geng et al., 2022). Consequently, LLMs may struggle to meet the nuanced needs of recommendation tasks. To alleviate these issues, one potential solution is to frequently fine-tune the LLMs with up-to-date and domain-specific knowledge. However, the massive parameters of LLMs make this process computationally expensive and time-consuming, which severely hinders the practical application in the real world.
@@
To address the aforementioned challenges, in this paper, we propose a knowledge retrieval-augmented recommendation framework, namely K-RagRec, to provide up-to-date and reliable knowledge by retrieving relevant knowledge from item’s KGs for recommendation generation in the era of LLMs. Specifically, our proposed framework first performs knowledge sub-graph indexing on the items’ KG at a coarse and fine granularity to construct the knowledge vector database. Next, a popularity selective retrieval policy is designed to determine which items should be retrieved, followed by the retrieval of specific sub-graphs from the knowledge vector database. To refine the quality of retrieval and ensure the most relevant results are prioritized at the top of the input, we subsequently re-rank the retrieved knowledge sub-graphs. Finally, we introduce a GNN and projector to align the retrieved knowledge sub-graphs into the semantic space of the LLM for knowledge-augmented recommendation. The main contributions of this paper can be summarized as follows:
  • We propose a novel framework that retrieves faithful knowledge from KGs to augment the recommendation capability of LLM. Note that we introduce a flexible indexing method for KGs, which can provide a comprehensive view of a node’s neighborhood within KG.
  • We design a popularity selective retrieval strategy to determine whether an item needs to be retrieved based on its popularity, significantly improving efficiency.
@@ 2 Related Work
Recently, RAG has emerged as one of the most representative technologies in the field of generative AI, combining the strengths of retrieval systems and language models (LM) to produce coherent and informative text. Early methods, such as REALM (Guu et al., 2020), RETRO (Borgeaud et al., 2022), and DPR (Karpukhin et al., 2020), typically involve retrieving relevant fragments from a large corpus to guide the LM generation. However, standard RAG methods often struggle to accurately retrieve all relevant textual chunks, due to unnecessary noise and even harmful disturbance in the documents. To address these limitations, recent studies (Baek et al., 2023; Wu et al., 2023b; He et al., 2024; Luo et al., 2023; Wang et al., 2023; Sen et al., 2023; Sun et al., 2023) focus on retrieving structured and faithful knowledge from graphs for enhancing generations. For example, Retrieve-Rewrite-Answer (Wu et al., 2023b) retrieves relevant sub-graphs from KG and converts retrieved sub-graphs into text for the generation. G-Retriever (He et al., 2024) explores retrieving sub-graphs from various types of graphs to alleviate the hallucinations of LLM.
With the explosion of LLMs in recommendations, a few works (Di Palma, 2023; Wu et al., 2024; Ang et al., 2024) make initial explorations in RAG for recommendations. For instance, the work (Di Palma, 2023) proposes leveraging knowledge from movie or book datasets to enhance recommendations. Nevertheless, retrieving faithful structured knowledge from the KGs for recommendations is under-explored and shows great potential. To fill this gap, we propose to retrieve knowledge sub-graphs from KGs to enhance the recommendation performance of LLM. We provide more comprehensive related work on Appendix A.3.
@@ 3.1 Preliminary
Knowledge Graph: In this work, we propose to leverage external knowledge databases (i.e., KGs) to augment LLMs for recommendations. KGs contain abundant factual knowledge in the form of a set of triples: $\mathcal{G}=\{(n,e,n^{\prime})\ |\ n,n^{\prime}\in N\ ,\ e\in E\}$, where $N$ and $E$ denote the set of entities and relations, respectively. For example, a triple $Interstellar\ \xrightarrow{directed\_by}\ Nolan$ indicates that the movie "Interstellar" was directed by the director "Nolan".
LLM-based Recommendations: Let $U=\{u_{1},u_{2},\dots,u_{n}\}$ and $V=\{v_{1},v_{2},\dots,v_{m}\}$ represent the sets of users and items, where $n$ and $m$ are the sizes of users and items, respectively. The goal of an LLM-based recommender system is to understand users’ preferences by modeling users’ historical items interactions $V^{u_{i}}=[v^{u_{i}}_{1},v^{u_{i}}_{2},\cdots,v^{u_{i}}_{|L_{u_{i}}|}]$ (e.g., clicks, bought, etc.), where $|L_{u_{i}}|$ is interaction sequence length for user $u_{i}$. Notably, item $v_{j}$’ side information, such as title and description, is publicly available to enhance LLM for modeling user preferences. In our setting, we consider asking the LLM to select user $u$’s preferred item $v$ from the candidate set $C=\{v_{j}\}^{M}_{j=1}$, where $M$ is the number of candidate items. The candidate set $C$ typically consists of one positive sample as well as $M-1$ negative samples. Specifically, for a frozen LLM $f_{\delta}$ with parameters $\delta$, we denote an input-output sequence pair as $(Q,A)$, where $Q$ is the recommendation query/prompt, which consists of task descriptions and users’ historical items. The output $A$ is the LLM’s prediction. Furthermore, we introduce the concept of GNNs in appendix A.2.
@@ 3.2 The Overview of Proposed Method
As shown in Figure 2, our proposed K-RagRec consists of five crucial components: Hop-Field Knowledge Sub-graphs for Semantic Indexing, Popularity Selective Retrieval Policy, Knowledge Sub-graphs Retrieval, Knowledge Sub-graphs Re-Ranking, and Knowledge-augmented Recommendation. The model first performs indexing of hop-field knowledge sub-graphs within the KG. Following this, a popularity selective retrieval policy is implemented to determine which items should be retrieved or augmented. The model then retrieves specific sub-graphs from the knowledge vector database. Subsequently, the retrieved knowledge sub-graphs are re-ranked to refine the retrieval quality. Finally, the retrieved knowledge sub-graphs are utilized with the original prompt to generate recommendations.
@@ 3.3 Hop-Field Knowledge Sub-graphs for Semantic Indexing
Figure 2: The overview of the K-RagRec. It contains five key components: Hop-Field Knowledge Sub-graphs for Semantic Indexing, Popularity Selective Retrieval Policy, Knowledge Sub-graphs Retrieval, Knowledge Sub-graphs Re-Ranking, and Knowledge-augmented Recommendation.
Typically, retrieving knowledge from KG chunks the KG into nodes (He et al., 2024) or triplets (Luo et al., 2023) and only retrieves content locally around the target entity in the KG. However, these methods just naively retrieve the first-order neighbors of an entity (i.e., item), which makes it difficult to capture these higher-order neighborhood effects among entities/items in the recommendation process. Therefore, to effectively retrieve knowledge from the KG, we propose performing semantic indexing on hop-field knowledge sub-graphs, which can flexibly chunk KGs and provide a comprehensive view of a node’s neighborhood in KG. As illustrated in Figure 2 component 1, we first introduce a pre-trained language model (PLM), such as SentenseBert (Reimers, 2019), to capture the semantic information for node $n_{o}$ as follows:
@@
$\textbf{z}_{n_{o}}^{l}=\text{GNN}_{\phi_{1}}^{\text{Indexing}}(\{\textbf{z}_{n_{m}}^{(l-1)},\textbf{r}_{e_{<o,m>}}^{(l-1)}:\\ ~{}n_{m}\in\mathcal{N}(n_{o})\}),$ (3)
@@
3.4 Popularity Selective Retrieval Policy
Although RAG can augment the LLM for modeling user preferences with retrieved knowledge, retrieving each item can cost a significant amount of retrieval time, which can severely degrade the user experience and cause user churn. Meanwhile, most users’ online behaviours in recommender systems are following the power law distribution (Abdollahpouri et al., 2017; Celma and Herrera, 2008) in which a small proportion (e.g., less than 20%) of items (i.e., popular items) often account for a large proportion of users’ online behaviours (e.g., more than 80%), while cold-start items have a few interactions from users. Therefore, most models tend to keep rich knowledge of popular items, resulting in an inferior performance for cold-start items (Zhao et al., 2023). To this end, we design a popularity selective retrieval policy to determine whether an item needs to be augmented from KG based on its popularity (e.g., sales volume and page view). Particularly, the item is retrieved if its popularity is less than the pre-defined threshold $p$, otherwise not. By incorporating this strategy, the retrieval time in K-RagRec can be significantly reduced to achieve more efficient retrieval.
@@ 3.5 Knowledge Sub-graphs Retrieval
Next, we retrieve the top-$K$ most similar sub-graphs $G_{j}=\{g^{\prime}_{1},...,g^{\prime}_{K}\}$ from $\mathcal{G}$ for item $v_{j}$:
@@ 3.6 Knowledge Sub-graphs Re-Ranking
To refine the quality of retrieval for enhancing recommendation performance, the next crucial step is knowledge sub-graphs re-ranking. Feeding all retrieved knowledge sub-graphs $g^{\prime}\in G$ directly into LLM can lead to information overload if the user ${u_{i}}$ has a long historical interaction list towards items. Therefore, we execute re-ranking to shorten the retrieved knowledge sub-graphs and ensure the most relevant knowledge sub-graphs are prioritized at the top of the prompt. Specifically, we adopt the recommendation prompt as a query for re-ranking, which consists of task descriptions and users’ historical items. For example, this recommendation prompt can be "What is the top recommended movie for the user who watched {Matrix, …, Iron Man}?". For consistency, we adopt the same PLM to capture the semantic information of the above prompt as p, and re-rank the knowledge sub-graphs in $G$ to obtain a Top-$N$ knowledge sub-graphs set $\hat{G}$:
@@ 3.8 Optimization for K-RagRec
Therefore, instead of fine-tuning the LLM model extensively, we only learn parameters of two GNNs (i.e., $\text{GNN}_{\phi_{1}}^{\text{Indexing}}$, $\text{GNN}_{\phi_{2}}^{\text{Encoding}}$) and projector $\text{MLP}_{\theta}$, while parameters $\delta$ of LLM backbone are frozen. We update the parameters $\phi_{1}$, $\phi_{2}$ and $\theta$ through the Cross-Entropy Loss $\mathcal{L}(Y,A)$, where $Y$ is the ground-truth and $A$ is LLM’s prediction.